katwhirles:

6/2(1+2)= whats this equal

CoolRekterizer:

What do you think?

katwhirles:

i need help answering it I forget how to do it

CoolRekterizer:

Well, 6/2 = 3 when you simplify it, and 1 + 2 = 3. Of course, you need to multiply both those numbers, and the answer would give you 9.

katwhirles:

thats what i was thinking to

CoolRekterizer:

@gracie101507 1 would not be the answer...

gracie101507:

how did you get 9

katwhirles:

but wouldnt you do 2 times 1 first

CoolRekterizer:

You got to put it like this, (6/2) x (1+2)

CoolRekterizer:

Do PEMDAS and you can get your answer.

Shadow:

\[6/2(1+2)= ?\] According to PEMDAS: Parentheses Exponents Multiplication Division Addition Subtraction We look to the parentheses first, simplifying to -> \[6/2(3)= ?\]| Since the only thing remaining is multiplication and division, we can do them in either order. Starting with division is 'prettier' though. \[3(3) = 9\]

CoolRekterizer:

Just like Shadow did.

katwhirles:

but wouldn't you do 2 times whats in the parenthesis first

CoolRekterizer:

Shadow is right, just look at the steps...

Shadow:

PEMDAS means you start what's INSIDE the parentheses first, not outside.

katwhirles:

ok thanks

Shadow:

No problem

katwhirles:

ok im still confused

Shadow:

With?

katwhirles:

that same question

gracie101507:

with that same question

Shadow:

What part of the question is confusing you?

gracie101507:

everything

gracie101507:

i still dont really know if it is 9

katwhirles:

the part where it says parenthesis first

katwhirles:

if you didn't do that you would have got 1

Shadow:

Nevermind I see the problem

katwhirles:

whats the problem

Shadow:

We were viewing it as a division function, which in reality, it's a fraction.

Shadow:

Actually wait, that doesn't make sense.

katwhirles:



Hero:

This is one of those "screenshot if from the book" questions

Shadow:

Lol

katwhirles:

no i just did that

Shadow:

Oh I see

Shadow:

You did the Distributive Property

katwhirles:

yea

Shadow:

I think the Distributive Property counts as multiplication, as opposed to being considered a parentheses function. Which is why you would do the addition inside of the parentheses first.

Shadow:

Since parentheses is ranked higher than multiplication in PEMDAS.

Hero:

I don't see where it was screenshot from the original text

Hero:

The way you wrote the original problem is ambiguous

Shadow:

Yeah if you can take a picture and upload it, or take a screenshot, that'd clear things up @katwhirles

katwhirles:

ok

Hero:

Since some will interpret it as \(\dfrac{6}{2}(1 + 2)\) and others will interpret it as \(\dfrac{6}{2(1 + 2)}\) which are two completely different expressions

Shadow:

Yeah but in order for it to be the second, the parentheses would have to include 2 and the 1 + 2, so that it would all be in the denominator, under 6. But it's just 6/2 multiplied by (1+2).

Shadow:

Of course, taking what @katwhirles wrote as what the textbook/problem depicts.

katwhirles:

gracie101507:

i think it is 1

gracie101507:

do u think it is 1?????